OpenStudy (anonymous):

If a given angle is 11pi/6, find A) Complement of the angle. B) Supplement of the angle.

OpenStudy (anonymous):

well 11pi/6 is 330 degrees.... so there really isn't a complement of the angle or a supplement of the angle... but i'm sure the answer would be 60 degrees (pi/3 radians) for the complement and 150 degrees (5pi/6 radians) for the supplement

OpenStudy (anonymous):

since 360 - 330 = 30 90 - 30 = 60 180 - 30 = 180

OpenStudy (anonymous):

OR you could think that 11pi/6 is -30 degrees... so the complement is 90 - (-30) = 120 degrees (2pi/3 radians) and the supplement is 180 - (-30) = 210 degrees (7pi/6 radians) it depends on how you look at it

OpenStudy (anonymous):

okay well isn't a complement suppose to add up to equal 90 ad a supplement 180?

OpenStudy (anonymous):

yes, it depends on what angle you consider 11pi/6

OpenStudy (anonymous):

oh alright i get it now! thanks for the help!!

OpenStudy (anonymous):

you're welcome!